OpenStudy (anonymous):

The perimeter of a rectangle is determined to be 5 times its width. Its semiperimeter (half of the perimeter) is calculated to be 10 inches. What is the area of the rectangle?

OpenStudy (ash2326):

perimeter=20 (semi is given 10) 2(l+b)=20 (l is the length and b is the breadth or width) also perimeter=5b 20=5b b=4 2(l+b)=20 l+b=10 l=10-b l=10-4 l=6 b=4

OpenStudy (anonymous):

Perimeter = 20 inches 5b=20 b=4 2(l+b)=20 l+b=10 l+4=10 l=6 Area = 6*4=24in^2

OpenStudy (ash2326):

sorry area=l*b area=6*4=24